# SafeFix: Targeted Model Repair via Controlled Image Generation

### v1

@@ Abstract
Deep learning models for visual recognition often exhibit systematic errors due to underrepresented semantic subpopulations. Although existing debugging frameworks can pinpoint these failures by identifying key failure attributes, repairing the model effectively remains difficult. Current solutions often rely on manually designed prompts to generate synthetic training images—an approach prone to distribution shift and semantic errors. To overcome these challenges, we introduce a model repair module that builds on an interpretable failure attribution pipeline. Our approach uses a conditional text-to-image model to generate semantically faithful and targeted images for failure cases. To preserve the quality and relevance of the generated samples, we further employ a large vision–language model (LVLM) to filter the outputs, enforcing alignment with the original data distribution and maintaining semantic consistency. By retraining vision models with this rare-case-augmented synthetic dataset, we significantly reduce errors associated with rare cases. Our experiments demonstrate that this targeted repair strategy improves model robustness without introducing new bugs. Code is available at https://github.com/oxu2/SafeFix.
@@ 1 Introduction
Despite extensive training on annotated images and evaluation on static test sets, computer vision models can still fail in unexpected ways Barbu et al. (2019); Gao et al. (2023); Leclerc et al. (2022). One common cause is dataset bias: certain subpopulations are underrepresented in the training data. For instance, facial recognition systems have shown significantly higher error rates for darker-skinned women compared to lighter-skinned men, due to imbalanced representation in training datasets Buolamwini and Gebru (2018). Therefore, identifying and repairing potential model failures caused by underrepresented semantic subpopulations before deployment is essential to improve the reliability of computer vision systems in real-world applications.
@@
Several recent studies have addressed the identification and repair of model failures caused by underrepresented semantic subpopulations. For example, given a natural language description of a potentially underrepresented semantic subpopulation, AdaVision Gao et al. (2023) introduces a model debugging pipeline that leverages CLIP Radford et al. (2021) to iteratively retrieve relevant images from LAION-5B Schuhmann et al. (2022) and evaluate model predictions. HiBug Chen et al. (2023) advances interpretable model debugging by automatically discovering failure slices based on shared visual attributes using a pre-trained vision–language model. DCD Singla et al. (2024) identifies failure modes by collecting misclassified samples, computing a failure direction that captures their distinguishing attributes, and retrieving similar unlabeled examples that reveal systematic model weaknesses and underrepresented subpopulations. The latent-space-based filtering method Jain et al. (2023b) identifies failure patterns by training a linear SVM in a vision–language latent space to separate correct from incorrect predictions, revealing interpretable subpopulation errors without requiring labels. In addition, D3M Jain et al. (2023a) proposes to remove harmful training samples that disproportionately hurt underrepresented groups, thus enhancing overall model performance.
@@
While existing methods can effectively identify model errors caused by underrepresented semantic subpopulations, the subsequent and critical step—model repair—still presents key challenges. For instance, AdaVision Gao et al. (2023) employs a simple fine-tuning strategy to retrain the model on retrieved images that cause model errors. However, this approach can degrade performance on images where the model originally performed well. HiBug Chen et al. (2023) addresses model repair by turning attributes describing certain underrepresented subpopulations into language prompts to synthesize new training images via generative models. However, the synthesized images may suffer from distribution mismatch compared to the original training data. Moreover, the repair process can introduce new bugs, as the generated images may fail to preserve attributes unrelated to the original model errors.
@@
Building on these observations, we propose SafeFix, a targeted model repair method designed to correct failures caused by underrepresented semantic subpopulations through controlled image generation—without degrading overall performance or introducing new errors. Compared to AdaVision Gao et al. (2023) and DCD Singla et al. (2024), which retrieve samples from another dataset, our method generates new examples grounded in the training set to ensure distributional consistency. Unlike HiBug Chen et al. (2023), which simply composes failure attributes into language prompts, SafeFix anchors generation on real instances in the training set to preserve unrelated attributes and avoid unintended changes. While the latent-space-based filtering method Jain et al. (2023b) explores repair via diffusion models, their approach relies solely on caption-based prompting and does not guarantee the semantic accuracy of critical attributes. In contrast, SafeFix mitigates the unreliability of generative models in rendering sensitive attributes—such as skin tone or hair color—by further filtering out semantically inconsistent images via a large vision–language model (LVLM) Bai et al. (2025); Liu et al. (2023). This ensures that the final outputs faithfully reflect the intended attribute change (e.g., generating a sad expression on a darker-skinned woman with red hair, whereas standard diffusion models often mistakenly produce a darker-skinned man with black hair instead), enabling precise and robust repair of underrepresented semantic subpopulation failures.
@@
SafeFix directly addresses two major challenges in existing model repair strategies: 1) ensuring the generated images reflect the true failure attributes, and 2) aligning them with the original data distribution. By retraining models on this augmentation set, we reduce errors associated with underrepresented semantic subpopulations, where model failures stem from insufficient training coverage of specific attributes.
Our contributions are as follows:
@@ 2 Related Work
Failure Pattern Discovery. HiBug Chen et al. (2023) proposes a pioneering pipeline that identifies interpretable failure cases in vision models by clustering semantically meaningful attributes, revealing both rare categories and spurious correlations. HiBug2 Chen et al. (2025) extends this approach with more efficient error-slice discovery and a closed-loop debugging mechanism, improving coherence and coverage of discovered model bugs. Several follow-up works attempt to strengthen interpretability in model debugging Adebayo et al. (2020, 2022), though often limited by static failure patterns or inadequate visual grounding. Beyond HiBug, MODE Vendrow et al. (2023) introduces a state-differential analysis framework that locates internal model faults and proposes data-driven remedies. TCAV Kim et al. (2018) further enriches interpretability by testing the model’s sensitivity to high-level concepts and by correcting spurious activations at the concept level. 3DB Leclerc et al. (2022) complements these efforts by constructing structured attribute spaces over failure modes, allowing discovery of underrepresented attributes through unsupervised analysis of visual model errors.
@@
Targeted Synthetic Augmentation via Diffusion Models. Diffusion models such as Stable Diffusion Rombach et al. (2022) and classifier-free guidance Ho and Salimans (2022) enable controllable, semantically faithful image synthesis. These models have proven useful in debugging Casper et al. (2022); Fang et al. (2024); Huang et al. (2024), counterfactual visualization Augustin et al. (2022); Boreiko et al. (2022), and training data augmentation Trabucco et al. (2023); Dunlap et al. (2023), especially in few-shot and fine-grained recognition tasks. Recent work like DiGA Zhang et al. (2024) shows how editing spurious attributes while preserving class semantics can mitigate bias without requiring new annotations. Others, such as GenMix Lee et al. (2024), propose mixing synthetic and real images to enhance robustness. These advances highlight how targeted generation can shape training distributions to address model weaknesses. Building on these insights, our method forms a closed-loop debugging pipeline that not only identifies failure cases but also synthesizes and integrates targeted images to improve model performance. Compared to prior augmentation or error discovery pipelines Fang et al. (2024); Huang et al. (2024); Chen et al. (2023), our approach is more generalizable and less reliant on predefined attribute sets. We draw inspiration from efforts like StylizedImageNet Geirhos et al. (2019), debiasing pipelines Jin and Rinard (2021), and adaptive augmentation methods Mikołajczyk-Bareła et al. (2023); Zhao et al. (2022); Wang et al. (2024b), but focus on semantically controllable generation tailored to discovered bugs.
@@
Multimodal Filtering via LVLMs. LVLMs like Flamingo Alayrac et al. (2022) and OpenFlamingo Awadalla et al. (2023) have demonstrated strong capabilities in semantically grounding visual concepts. Recent studies have explored using LVLMs themselves as filters to select high-quality image-text pairs, enhancing dataset quality for downstream tasks Wang et al. (2024a); Li et al. (2024). These approaches outperform traditional methods like CLIP-based filtering by providing fine-grained, attribute-aware analysis of generated samples. In line with these trends, we adopt an LVLM as an automated filtering component, grounding our approach in established methods that leverage LVLMs for semantic validation of generated data.
@@
Targeted Repair for Rare-Case Bugs. Recent work explores how underrepresented subpopulations induce systematic errors in vision models and how targeted interventions can mitigate such failure modes. DOMINO Eyuboglu et al. (2022) discovers coherent failure slices by clustering model errors in a cross-modal embedding space, enabling automated identification of rare-case bugs without manual slice definitions. REAL Parashar et al. (2024) focuses on rare visual concepts that are neglected in large-scale vision–language datasets, augmenting them by retrieving semantically similar examples and fine-tuning lightweight classifiers on the retrieved subsets to improve model robustness on these rare categories. This work reflects a broader shift toward targeted model repair using interpretable diagnostics and subpopulation-aware interventions, aligning with our controlled synthesis and refinement approach.
@@ 3 Background
Let $x\in\mathcal{X}$ be an input image with its corresponding ground-truth label $y(x)\in\mathcal{Y}$. A computer vision model $f_{\theta}\colon\mathcal{X}\to\mathcal{Y}$ produces a prediction $f_{\theta}(x)$ for each input image $x$. Denote the training and validation splits by $\mathcal{D}_{\mathrm{train}}$ and $\mathcal{D}_{\mathrm{val}}$, respectively. The full dataset is then given by $\mathcal{D}=\mathcal{D}_{\mathrm{train}}\cup\mathcal{D}_{\mathrm{val}}$. The overall validation accuracy is computed as:
@@
In attribute-based model debugging Chen et al. (2023), each image is assigned several attributes that represent different subpopulations. In particular, let $\mathcal{A}=\{a_{1},\dots,a_{m}\}$ be a set of attribute functions, where each $a_{i}\colon\mathcal{X}\to\mathcal{V}_{i}$ maps an image $x$ to a discrete value $v\in\mathcal{V}_{i}$ and $\mathcal{V}_{i}$ represents the set of all possible values for attribute $i$. A slice $S$ is defined as the set of images that satisfy a conjunction of attribute-value assignments:
@@
$\mathrm{Acc}(S)=\frac{1}{|S|}\sum_{x\in S}\mathbf{1}\{f_{\theta}(x)=y(x)\}.$ (3)
@@
Rare-case bugs, often caused by underrepresented semantic subpopulations Eyuboglu et al. (2022), are attribute-based failure modes for which the model’s error rate on validation samples matching a target description significantly exceeds its overall error and those samples appear infrequently in the training dataset. We say a slice $S_{r}$ is a rare-case slice if it constitutes less than a fraction $\rho$ of the training data, i.e.,
@@ 4 SafeFix
We begin by generating visually controlled images using a text-to-image diffusion model, Stable Diffusion, guided by a structure-conditioned controller, ControlNet Zhang et al. (2023). Instead of relying solely on language prompts, we generate images conditioned on training data while modifying specific attributes to reflect failure slice semantics. Next, we employ an LVLM to automatically verify whether each generated image accurately reflects the intended attributes. This filtering step ensures that only semantically faithful samples are retained. Finally, we augment the original training dataset with the validated images and retrain the model. SafeFix enhances performance on error-prone regions without compromising overall accuracy or introducing new bugs.
@@ 4.1 Model Diagnosis for Identifying Rare-Case Bugs
We begin by training a standard computer vision model $f_{\theta}$ on the original training set $\mathcal{D}_{\mathrm{train}}$. After obtaining predictions on the validation set $\mathcal{D}_{\mathrm{val}}$, similar to HiBug, we use a large vision-language model (LVLM; e.g., GPT-4 with vision OpenAI (2023)) to propose candidate attributes and a VQA model (e.g., BLIP Li et al. (2022)) to assign attribute values $a_{i}(x)$ to each image. We extract a set of rare-case bug slices $\{S_{e}\}$, each defined by a conjunction of attribute conditions that exhibit both high error rates and low coverage in the dataset. Specifically, to identify underrepresented and error-prone subpopulations, we analyze each attribute $a_{i}$ and its associated values $v\in\mathcal{V}_{i}$. For each value $v$, we examine the slice $S=\{x\mid a_{i}(x)=v\}$ and compute two quantities: (1) its proportion in the validation set, and (2) the model’s accuracy on the corresponding validation samples.
We flag a value $v$ as a rare-case bug if:
  • The validation support $|S\cap\mathcal{D}_{\mathrm{val}}|$ is below the threshold $\rho\cdot|\mathcal{D}_{\mathrm{val}}|$.
  • The validation accuracy $\mathrm{Acc}(S)$ is significantly below the overall validation accuracy (by a fixed margin $\epsilon$).
@@ 4.2 Targeted Generation with Conditional Diffusion Models (CDMs)
Next, we aim to produce attribute-preserving edits on each original image $x$, focusing on targeted attributes identified in problematic slices from the previous diagnostic stage. Specifically, we generate visually controlled synthetic images $x^{\prime}$ using a text-to-image diffusion model, Stable Diffusion, guided by the structure-conditioned controller ControlNet Zhang et al. (2023).

### v2

@@ Abstract
Deep learning models for visual recognition often exhibit systematic errors due to underrepresented semantic subpopulations. Although existing debugging frameworks can pinpoint these failures by identifying key failure attributes, repairing the model effectively remains difficult. Current solutions often rely on manually designed prompts to generate synthetic training images—an approach prone to distribution shift and semantic errors. To overcome these challenges, we introduce a model repair module that builds on an interpretable failure attribution pipeline. Our approach uses a conditional diffusion text-to-image model to generate semantically faithful and targeted images for failure cases. To preserve the quality and relevance of the generated samples, we further employ a large vision–language model (LVLM) to filter the outputs, enforcing alignment with the original data distribution and maintaining semantic consistency. The effectiveness of this filtering step is verified by human audit. By retraining vision models with the synthetic dataset and source dataset, we significantly reduce errors associated with rare cases. Our experiments demonstrate that this targeted repair strategy improves model robustness without introducing new bugs. Code is available at https://github.com/oxu2/SafeFix.
@@ 1 Introduction
Despite extensive training on annotated images and evaluation on static test sets, computer vision models can still fail in unexpected ways [barbu2019objectnet, gao2023adaptive, leclerc20223db]. One common cause is dataset bias: certain subpopulations are underrepresented in the training data. For instance, facial recognition systems have shown significantly higher error rates for darker-skinned women compared to lighter-skinned men, due to imbalanced representation in training datasets [buolamwini2018gender]. Therefore, identifying and repairing potential model failures caused by underrepresented semantic subpopulations before deployment is essential to improve the reliability of computer vision systems in real-world applications.
@@
Several recent studies have addressed the identification and repair of model failures caused by underrepresented semantic subpopulations. For example, given a natural language description of a potentially underrepresented semantic subpopulation, AdaVision [gao2023adaptive] introduces a model debugging pipeline that leverages CLIP [radford2021learning] to iteratively retrieve relevant images from LAION-5B [schuhmann2022laion] and evaluate model predictions. HiBug [chen2023hibug] advances interpretable model debugging by automatically discovering failure slices based on shared visual attributes using a pre-trained vision–language model. DCD [singla2024data] identifies failure modes by collecting misclassified samples, computing a failure direction for distinguishing attributes, and retrieving similar unlabeled examples to expose systematic weaknesses and underrepresented subpopulations. The latent-space method [jain2023latentfailures] discovers failure patterns by training a linear SVM in a latent space, separating correct from incorrect predictions to yield subpopulation errors without labels. D3M [jain2023d3m] enhances overall model performance by removing harmful training samples that disproportionately affect underrepresented groups.
@@
While existing methods can effectively identify model errors caused by underrepresented semantic subpopulations, the subsequent and critical step—model repair—still presents challenges. For instance, AdaVision [gao2023adaptive] employs a simple fine-tuning strategy to retrain the model on retrieved error-causing images. However, this approach can degrade performance on images where the model originally performed well. HiBug [chen2023hibug] addresses model repair by turning attributes describing underrepresented subpopulations into language prompts to synthesize new training images via generative models. However, the synthesized images may suffer from distribution mismatch compared to the original training data. Moreover, the repair process can introduce new bugs, as the generated images may fail to preserve attributes unrelated to original model errors.
@@
Building on these observations, we propose SafeFix, a targeted model repair method designed to correct failures caused by underrepresented semantic subpopulations through controlled image generation—without degrading overall performance or introducing new errors. Compared to AdaVision [gao2023adaptive] and DCD [singla2024data], which retrieve samples from another dataset, our method generates new examples grounded in the training set to ensure distributional consistency. Unlike HiBug [chen2023hibug], which simply composes failure attributes into language prompts, SafeFix anchors generation on real instances in the training set to preserve unrelated attributes and avoid unintended changes. While the latent-space-based filtering method [jain2023latentfailures] explores repair via diffusion models, their approach relies solely on caption-based prompting and does not guarantee the semantic accuracy of critical attributes. In contrast, SafeFix mitigates the unreliability of generative models in rendering sensitive attributes—such as skin tone or hair color—by further filtering out semantically inconsistent images using a large vision–language model (LVLM) [bai2025qwen25vl, liu2023improvedllava], and the LVLM is verified by human audit. This ensures that the final outputs faithfully reflect the intended attribute change (e.g., generating a sad expression on a darker-skinned woman with red hair, whereas standard diffusion models often mistakenly produce a darker-skinned man with black hair instead), enabling precise and robust repair of underrepresented semantic subpopulation failures.
@@
SafeFix directly addresses two major challenges in existing model repair strategies: 1) ensuring the generated images reflect the true failure attributes, and 2) aligning them with the original data distribution. By retraining models on this augmentation set, we reduce errors associated with underrepresented semantic subpopulations, where model failures stem from insufficient training coverage of specific attributes. Our contributions are as follows:
@@ 2 Related Work
Failure Pattern Discovery. HiBug [chen2023hibug] proposes a pioneering pipeline that identifies interpretable failure cases in vision models by clustering semantically meaningful attributes, revealing both rare categories and spurious correlations. HiBug2 [chen2025hibug2] extends this approach with more efficient error-slice discovery and a closed-loop debugging mechanism, improving coherence and coverage of discovered model bugs. Several follow-up works attempt to strengthen interpretability in model debugging [adebayo2020debugging, adebayo2022posthoc], though often limited by static failure patterns or inadequate visual grounding. Beyond HiBug, MODE [vendrow2023dataset] introduces a state-differential analysis framework that locates internal model faults and proposes data-driven remedies. TCAV [kim2018tcav] further enriches interpretability by testing the model’s sensitivity to high-level concepts and by correcting spurious activations at the concept level. 3DB [leclerc20223db] complements these efforts by constructing structured attribute spaces over failure modes, allowing discovery of underrepresented attributes through unsupervised analysis of visual model errors.
@@
Targeted Synthetic Augmentation via Diffusion Models. Diffusion models such as Stable Diffusion [rombach2022high] and classifier-free guidance [ho2022classifierfree] enable controllable, semantically faithful image synthesis. These models have proven useful in debugging [casper2022robust, fang2024data, huang2024data], counterfactual visualization [augustin2022diffusion, boreiko2022sparse], and training data augmentation [trabucco2023effective, dunlap2023diversify], especially in few-shot and fine-grained recognition tasks. Recent work like DiGA [zhang2024distributionally] shows how editing spurious attributes while preserving class semantics can mitigate bias without requiring new annotations. Others, such as GenMix [lee2024genmix], propose mixing synthetic and real images to enhance robustness. These advances highlight how targeted generation can shape training distributions to address model weaknesses. Building on these insights, our method forms a closed-loop debugging pipeline that not only identifies failure cases but also synthesizes and integrates targeted images to improve model performance. Compared to prior augmentation or error discovery pipelines [fang2024data, huang2024data, chen2023hibug], our approach is more generalizable and less reliant on predefined attribute sets. We draw inspiration from efforts like StylizedImageNet [geirhos2019imagenet], debiasing pipelines [jin2021background], and adaptive augmentation methods [mikolajczyk2023targeted, zhao2022adaptive, wang2024domain], but focus on semantically controllable generation tailored to discovered bugs.
@@
Multimodal Filtering via LVLMs. LVLMs like Flamingo [alayrac2022flamingo] have demonstrated strong capabilities in semantically grounding visual concepts. Recent studies show using LVLMs as filters to select high-quality image-text pairs enhanced dataset quality for downstream tasks [wang2024finetuned, li2024self]. These approaches outperform traditional methods like CLIP-based filtering by providing fine-grained, attribute-aware analysis of generated samples. In line with these trends, we adopt an LVLM as an automated filtering component, grounding our approach in established methods that leverage LVLMs for semantic validation of generated data.
@@
Targeted Repair for Rare-Case Bugs. Recent work explores how underrepresented subpopulations induce systematic errors in vision models and how targeted interventions can mitigate such failure modes. DOMINO [eyuboglu2022rare] discovers coherent failure slices by clustering model errors in a cross‑modal embedding space, enabling automated identification of rare-case bugs without manual slice definitions. REAL [parashar2024neglected] focuses on rare visual concepts that are neglected in large-scale vision–language datasets, augmenting them by retrieving semantically similar examples and fine-tuning lightweight classifiers on the retrieved subsets to improve model robustness on these rare categories. This work reflects a broader shift toward targeted model repair using interpretable diagnostics and subpopulation-aware interventions, aligning with our controlled synthesis and refinement approach.
@@ 3 Background
Let $x\in\mathcal{X}$ be an input image with its corresponding ground-truth label $y(x)\in\mathcal{Y}$. A computer vision model $f_{\theta}\colon\mathcal{X}\to\mathcal{Y}$ produces a prediction $f_{\theta}(x)$ for each input image $x$. Denote the training, validation, and test splits by $\mathcal{D}_{\mathrm{train}}$, $\mathcal{D}_{\mathrm{val}}$, and $\mathcal{D}_{\mathrm{test}}$, respectively. The full dataset is then given by $\mathcal{D}=\mathcal{D}_{\mathrm{train}}\cup\mathcal{D}_{\mathrm{val}}\cup\mathcal{D}_{\mathrm{test}}$. The overall validation accuracy is computed as:
@@
In attribute-based model debugging [chen2023hibug], each image is assigned several attributes that represent different subpopulations. In particular, let $\mathcal{A}=\{a_{1},\dots,a_{m}\}$ be a set of attribute functions, where each $a_{i}\colon\mathcal{X}\to\mathcal{V}_{i}$ maps an image $x$ to a discrete value $v\in\mathcal{V}_{i}$ and $\mathcal{V}_{i}$ represents the set of all possible values for attribute $i$. A slice $S$ is defined as the set of images that satisfy a conjunction of attribute-value assignments:
@@
$\mathrm{Acc}(S)=\frac{1}{|S|}\sum_{x\in S}\mathbf{1}\{f_{\theta}(x)=y(x)\}.\vskip-5.0pt$ (3)
@@
Rare-case bugs, often caused by underrepresented semantic subpopulations [eyuboglu2022rare], are attribute-based failure modes for which the model’s error rate on validation samples matching a target description significantly exceeds its overall error and those samples appear infrequently in the training dataset. We say a slice $S_{r}$ is a rare-case slice if it constitutes less than a fraction $\rho$ of the training data, i.e.,
@@ 4 SafeFix
We begin by generating visually controlled images using a text-to-image diffusion model, Stable Diffusion, guided by a structure-conditioned controller, ControlNet [zhang2023adding]. Instead of relying solely on language prompts, we generate images conditioned on training data while modifying specific attributes to reflect failure slice semantics. Next, we employ an LVLM to automatically verify whether each generated image accurately reflects the intended attributes. This filtering step ensures that only semantically faithful samples are retained. Finally, we augment the original training dataset with the validated images and retrain the model. SafeFix enhances performance on error-prone regions without compromising overall accuracy or introducing new bugs.
@@ 4.1 Model Diagnosis for Identifying Rare-Case Bugs
We begin by training a standard computer vision model $f_{\theta}$ on the original training set $\mathcal{D}_{\mathrm{train}}$. After obtaining predictions on the validation set $\mathcal{D}_{\mathrm{val}}$, similar to HiBug, we use a large vision-language model (LVLM; e.g., GPT-4 with vision [openai2023gpt4]) to propose candidate attributes and a VQA model (e.g., BLIP [li2022blip]) to assign attribute values $a_{i}(x)$ to each image. We extract a set of rare-case bug slices $\{S_{e}\}$, each defined by a conjunction of attribute conditions that exhibit both high error rates and low coverage in the dataset. Specifically, to identify underrepresented and error-prone subpopulations, we analyze each attribute $a_{i}$ and its associated values $v\in\mathcal{V}_{i}$. For each value $v$, we examine the slice $S=\{x\mid a_{i}(x)=v\}$ and compute two quantities: (1) its proportion in the training set, and (2) the model’s accuracy on the corresponding validation samples.
We flag the slice $S$ as a rare-case bug if:
  • The training support $|S\cap\mathcal{D}_{\mathrm{train}}|$ is below the threshold $\rho\cdot|\mathcal{D}_{\mathrm{train}}|$.
  • The validation accuracy $\mathrm{Acc}(S)$ is significantly below the overall validation accuracy $\mathrm{Acc}(\mathcal{D}_{\mathrm{val}})$ (by a margin $\epsilon$).
@@ 4.2 Targeted Generation with Conditional Diffusion Models (CDMs)
Next, we aim to produce attribute-preserving edits on each original image $x$, focusing on targeted attributes identified in problematic slices from the previous diagnostic stage. Specifically, we generate visually controlled synthetic images $x^{\prime}$ using a text-to-image diffusion model, Stable Diffusion, guided by the structure-conditioned controller ControlNet [zhang2023adding].